After 6 games played in a competition, Mavericks Football Club averaged 1.5 goals per game. How many goals would they have to score in the seventh game for that average to increase to 2 goals per game?

1
Expert's answer
2022-01-10T17:45:33-0500

"\frac{1.5*6+x}{7}=2."

"9+x=14."

"x=5."

They have to score 5 goals in the seventh game.